# Demographic-aware fine-grained classification of pediatric wrist fractures

### v5

@@ Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate. This study addresses the problem using a multifaceted approach: framing it as a fine-grained recognition task, fusing patient metadata with X-rays, and leveraging weights from a separate fine-grained dataset rather than from a coarse-grained dataset like ImageNet. Unlike prior work, this is the first application of metadata integration for wrist pathology recognition. Our results show that our approach improves diagnostic accuracy by 2% on small custom curated dataset and over 10% on a larger fracture dataset.
Index Terms—  Pediatric fracture classification, fracture detection, medical X-ray imaging, fine-grained classification, Patient metadata
@@ 1 Introduction
Wrist fractures represent up to 75% of hand injuries in emergency departments and pose significant risk of long-term impairment if misdiagnosed [1]. Given the wrist’s complex anatomy and developmental variability, accurate radiographic interpretation is essential. Recent advances in computer vision have demonstrated promise for automated abnormality detection in musculoskeletal imaging. However, such methods typically rely on large, labeled datasets, which are scarce in medical domains, and often process X-rays without contextual patient information.
To overcome the limitations of image-only learning, multimodal frameworks that fuse patient metadata with imaging have been successfully applied in dermatology and oncology. Studies have shown consistent accuracy improvements by incorporating age, sex, and anatomical site information alongside dermoscopic or histopathology images [2, 3, 4, 5, 6]. For example, Thomas et al. [6] fused demographic variables with dermoscopy images, while Cai et al. [4] applied ViTs with metadata encoders, achieving 93.8% melanoma classification accuracy. Recently, Hinterwimmer et al. [7] extended this approach to musculoskeletal imaging, reporting gains in bone tumour classification by integrating X-rays and metadata.
@@
Despite these advances, metadata integration has not been investigated for wrist pathology recognition. This gap is significant, as age- and sex-specific skeletal variations critically influence wrist imaging. Carpal ossification follows predictable stages, growth plate closure occurs at different ages across sexes [8, 9], and pediatric fracture types differ from adult patterns [10]. Without demographic context, models may misclassify normal developmental features, such as unfused growth plates, as abnormalities.
We address this gap by (1) introducing the first framework that fuses radiographs with demographic metadata, (2) formulating the task as fine-grained recognition using transformer-based modeling, and (3) showing that fine-grained pretraining further boosts generalization rather than pretraining on ImageNet.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [11, 12]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Fig. 2 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
Table 1: Instances in training, validation, and test sets.
Three Wrist Pathologies (Limited set) Fracture vs. No Fracture (Entire set)
Class Training Validation Test Class Training Validation Test
Boneanomaly 392 98 119 Fracture 11854 1357 338
Fracture 400 100 120 No Fracture 10164 1357 338
@@
Total 1168 292 354 Total 22018 2714 676
@@ 2.2 Architectural Details
Downsampling is applied at the beginning of each stage. In S3 and S4, overlapping patch embedding is implemented using convolutions with a stride of 2 and zero-padding, effectively reducing resolution while maintaining contextual overlap [15]. Table 2 outlines the configuration details of the three variants of the architecture used in our experiments.
@@
$\textit{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}+B\right)V$ (1)
Here, $Q$, $K$, and $V$ are the query, key, and value matrices of size $\mathbb{R}^{((M^{2}+N)\times d)}$, where $M^{2}$ corresponds to the number of image patches and $N$ is the count of extra tokens (e.g., class and meta tokens). The bias matrix $\hat{B}$ is parameterized as described in [16], supporting relative positions from M1-M-1 to $M+1$. A shared bias is applied for non-visual tokens since their spatial position is undefined.
The Relative Transformer Block comprises multi-head self-attention (MSA), layer normalization (LN), and a multi-layer perceptron (MLP), applied to the input token sequence $z_{0}$:
$z_{0}$ $=[x_{\text{class}};x^{1}_{\text{meta}},\ldots,x^{n-1}_{\text{meta}};x^{1}_{\text{vision}},\ldots,x^{m}_{\text{vision}}]$ (2)
$z^{\prime}_{i}$ $=\text{MSA}(\text{LN}(z_{i-1}))+z_{i}$
$z_{i}$ $=\text{MLP}(\text{LN}(z^{\prime}_{i}))+z^{\prime}_{i}$
In this formulation, $z_{i}\in\mathbb{R}^{(M^{2}+N)\times d}$ captures the updated token sequence at each transformer layer.
Stages S3 and S4 each generate a class token, $z^{1}_{\text{class}}$ and $z^{2}_{\text{class}}$, which encode combined visual and meta representations. Since these tokens may differ in dimension, $z^{1}_{\text{class}}$ is first projected to match the shape of $z^{2}_{\text{class}}$ using an MLP. The two are then concatenated and passed through a 1D convolution followed by normalization:
$\hat{z}^{\text{1}}_{\text{class}}$ $=\text{MLP}(\text{LN}(z^{\text{1}}_{\text{class}}))$ (3)
$z_{\text{class}}$ $=\text{Conv1d}(\text{Concat}(\hat{z}^{\text{1}}_{\text{class}},z^{\text{2}}_{\text{class}}))$
$y$ $=\text{LN}(z_{\text{class}})$
The final output $y$ aggregates information across multiple scales and token types. When using external attributes as metadata, these are encoded as vectors and embedded using a fully connected non-linear function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$.
To maintain the model’s ability to learn important visual features, especially when metadata is less informative or potentially overfitting, a regularization strategy is applied during training that progressively masks parts of the meta-information. This balances the reliance on both vision and auxiliary data.
@@ 2.3 Experimental Settings
We compare against common CNNs pre-trained on ImageNet [17, 18, 19, 20, 21, 22, 23, 24, 25], including ViT, VGG16, GoogleNet, EfficientNetV2, AlexNet, DenseNet201, InceptionV3, and ResNet50. All models underwent training for 100 epochs using a resolution of 224x224 and a batch size of 32. Adam optimizer was utilized with a learning rate $10^{-3}$. For MetaFormer, however, the base learning rate was $10^{-4}$.
@@ 3 Results & Discussion
Table 3: Performance evaluation of different baseline neural networks and the MetaFormer on only image data.
Model Test Accuracy (Vision Only)
ViT 62.7%
VGG16 63.3%
GoogleNet 70.3%
EfficientNetV2 72.3%
AlexNet 74.3%
DenseNet201 76.8%
InceptionV3 77.4%
ResNet50 77.7%
MetaFormer FGVR (Base) 79.4%
Table 4: Accuracy of different MetaFormer configurations on vision, vision + meta through early fusion, and late fusion.
Config Vision Early Fusion Late Fusion
FG-0 79.4% 79.1% 80.5%
FG-1 78.5% 80.5% 79.1%
FG-2 77.4% 78.5% 79.4%
FG-2-inat (Ours) 79.9% 81.1% 81.4%
Table 5: Evaluation of our configuration of MetaFormer (FG-2-inat) using sensitivity, specificity, and precision.
Configuration Cls Sensitivity Specificity Precision
Vision Only 0 62.2% 98.7% 96.1%
1 100.0% 82.9% 75.0%
2 77.4% 88.3% 76.1%
Vision+Meta 0 66.4% 98.7% 96.3%
1 100.0% 82.9% 75.0%
2 77.4% 90.4% 79.5%
@@ 3.1 Baseline Comparison
Table 3 shows that MetaFormer outperforms widely CNNs. This gives credence to the fact that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal.
3.2 Metadata Integration & Pretraining
@@
Table 6: Accuracies obtained by our configuration of MetaFormer on individual metadata attributes and when both attributes are considered.
Attribute Early Fusion Late Fusion
Vision + Age 80.5% 79.4%
Vision + Gender 79.9% 80.8%
Vision + Age + Gender 81.1% 81.4%
Table 7: Performance evaluation on significantly larger (fracture vs. no fracture) dataset.
FG-2-inat (Ours) Test Accuracy
Vision Only 50.1%
Vision + Meta 60.4%
We then selected the “FG-2” configuration and used the weights obtained from training it on the iNaturalist dataset [26] to create our “FG-2-inat” configuration. This surpassed all other configurations, whether using image data alone or in conjunction with metadata. This configuration reached 81.4%, a 1.5% improvement over image-only training. It seems that the fine-grained pretraining improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate. Table 5 shows per-class sensitivity, specificity, & precision scores for configuration “FG-2-inat”. Similarly, as shown in Fig. 4, metadata enhances fracture detection while preserving performance in bone anomaly and soft tissue recognition.
@@
3.3 Interpretability
GradCAM heatmaps (Fig. 3) demonstrate that MetaFormer attends to clinically relevant wrist regions, particularly around fracture lines and anomalous bone structures. While most heatmaps align well with pathology, some exhibit diffuse attention, suggesting occasional reliance on spurious features. This indicates that while metadata integration improves classification, further refinement is needed for precise localization. Nevertheless, the alignment of salient regions with clinical expectations strengthens confidence in the model’s decision process.
@@ 4 Conclusion
Overall, our experiments demonstrate three key findings: (1) Fine-grained transformer architecture outperform conventional CNNs for pediatric wrist pathology. (2) Demographic metadata (age, sex) significantly improves recognition across both limited and large-scale datasets, with late fusion being most effective. (3) Fine-grained pretraining enhances generalization compared to using ImageNet weights. These results emphasize that integrating developmental context into vision models is essential for robust pediatric musculoskeletal diagnosis.

### v6

@@ Abstract
Wrist pathology recognition from radiographs is challenging because normal appearance varies markedly with age and sex. In pediatric imaging, evolving carpal ossification and open growth plates can resemble fractures, while sex-dependent timing of physeal closure changes key visual cues. Image-only models therefore risk confusing developmental anatomy for pathology, especially in small medical datasets. We address this by framing wrist diagnosis as a fine-grained visual recognition (FGVR) task and introducing a multimodal transformer that fuses X-rays with demographic metadata (age and sex). Unlike prior work, this is the first study to integrate metadata for wrist pathology recognition, and we further show that fine-grained pretraining transfers better than coarse ImageNet initialization. Our approach improves accuracy by 2% on a small curated dataset and by over 10% on a larger fracture dataset.
Index Terms—  Pediatric fractures, fracture detection, X-rays, fine-grained classification, Patient metadata
@@ 1 Introduction
Wrist fractures represent up to 75% of hand injuries in emergency departments and pose a significant risk of long-term impairment if misdiagnosed [8]. Given the wrist’s complex anatomy and developmental variability, accurate radiographic interpretation is essential. Recent advances in computer vision have demonstrated promise for automated abnormality detection in musculoskeletal imaging. However, such methods typically rely on large, labeled datasets, which are scarce in medical domains, and often process X-rays without contextual patient information.
To overcome the limitations of image-only learning, multimodal frameworks that fuse patient metadata with imaging have been successfully applied in dermatology and oncology. Studies have shown consistent accuracy improvements by incorporating age, sex, and anatomical site information alongside dermoscopic or histopathology images For example, Thomas et al. [18] fused demographic variables with dermoscopy images, while Cai et al. [2] applied ViTs with metadata encoders, achieving 93.8% melanoma classification accuracy. Recently, Hinterwimmer et al. [7] extended this approach to musculoskeletal imaging, reporting gains in bone tumour classification by integrating X-rays and metadata.
@@
Despite these advances, metadata integration has not been investigated for wrist pathology recognition. This gap is significant, as age- and sex-specific skeletal variations critically influence wrist imaging. Carpal ossification follows predictable stages, growth plate closure occurs at different ages across sexes [13, 21], and pediatric fracture types differ from adult patterns [rosendahl_myklebust_ulriksen_nøttveit_eide_djuve_brudvik_2021]. Without demographic context, models may misclassify normal developmental features, such as unfused growth plates, as abnormalities.
We address this gap by (1) introducing the first framework that fuses radiographs with demographic metadata, (2) formulating the task as FGVR using transformer-based modeling, and (3) showing that fine-grained pretraining further boosts generalization rather than pretraining on ImageNet.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [12]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [1]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Fig. 2 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
Table 1: Instances in training, validation, and test sets from the curated datasets.
Three Wrist Pathologies (Limited set)
Class Training Validation Test
Boneanomaly 392 98 119
Fracture 400 100 120
@@
Total 1168 292 354
Fracture vs. No Fracture (Entire set)
Class Training Validation Test
Fracture 11854 1357 338
No Fracture 10164 1357 338
Total 22018 2714 676
@@ 2.2 Architectural Details
Downsampling is applied at the beginning of each stage. In S3 and S4, overlapping patch embedding is implemented using convolutions with a stride of 2 and zero-padding, effectively reducing resolution while maintaining contextual overlap. Table 2 outlines the configuration details of the three variants of the architecture used in our experiments.
@@
$\textit{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}+B\right)V.$ (1)
Here, $Q$, $K$, and $V$ are the query, key, and value matrices of size $\mathbb{R}^{((M^{2}+N)\times d)}$, where $M^{2}$ corresponds to the number of image patches and $N$ is the count of extra tokens (e.g., class and meta tokens). The bias matrix $\hat{B}$ is parameterized, supporting relative positions from M1-M-1 to $M+1$. A shared bias is applied for non-visual tokens since their spatial position is undefined. The Relative Transformer Block comprises multi-head self-attention (MSA), layer normalization (LN), and a multi-layer perceptron (MLP), applied to the input token sequence $z_{0}$:
$z_{0}$ $=[x_{\text{class}};x^{1}_{\text{meta}},\ldots,x^{n-1}_{\text{meta}};x^{1}_{\text{vision}},\ldots,x^{m}_{\text{vision}}],$ (2)
$z^{\prime}_{i}$ $=\text{MSA}(\text{LN}(z_{i-1}))+z_{i},$
$z_{i}$ $=\text{MLP}(\text{LN}(z^{\prime}_{i}))+z^{\prime}_{i}.$
In this formulation, $z_{i}\in\mathbb{R}^{(M^{2}+N)\times d}$ captures the updated token sequence at each transformer layer. Stages S3 and S4 each generate a class token, $z^{1}_{\text{class}}$ and $z^{2}_{\text{class}}$, which encode combined visual and meta representations. Since these tokens may differ in dimension, $z^{1}_{\text{class}}$ is first projected to match the shape of $z^{2}_{\text{class}}$ using an MLP. The two are then concatenated and passed through a 1D convolution followed by normalization:
$\hat{z}^{\text{1}}_{\text{class}}$ $=\text{MLP}(\text{LN}(z^{\text{1}}_{\text{class}})),$ (3)
$z_{\text{class}}$ $=\text{Conv1d}(\text{Concat}(\hat{z}^{\text{1}}_{\text{class}},z^{\text{2}}_{\text{class}})),$
$y$ $=\text{LN}(z_{\text{class}}).$
The final output $y$ aggregates information across multiple scales and token types. When using external attributes as metadata, these are encoded as vectors and embedded using a fully connected non-linear function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$. To prevent overreliance on metadata (and potential overfitting), we apply progressive metadata masking during training, forcing the model to learn robust visual features while still leveraging auxiliary signals.
@@ 2.3 Experimental Settings
We compare against common CNNs pre-trained on ImageNet [3, 5, 14, 15, 17, 10, 9, 16, 6, 20, 11], as shown in Table 3. All models underwent training for 100 epochs using a resolution of $224\times 224$ and a batch size of 32. Adam optimizer was utilized with $lr=10^{-3}$ for baselines and AdamW with base $lr=10^{-4}$ were defaults for MetaFormer.
@@ 3 Results & Discussion
Table 3: Vision-only baselines on the test set. MetaFormer fusion variants (bottom) achieve the best performance.
Model Accuracy (%) (95% CI)
ViT $59.0\ (54.0\text{--}64.1)$
AlexNet $66.1\ (61.3\text{--}71.2)$
YOLO11s $69.5\ (64.7\text{--}74.3)$
VGG16 $70.1\ (65.5\text{--}74.9)$
GoogleNet $72.9\ (68.1\text{--}77.4)$
ResNet50 $74.0\ (69.8\text{--}78.5)$
DenseNet201 $74.9\ (70.3\text{--}79.4)$
ConvNeXtV2-Tiny $76.0\ (71.5\text{--}80.2)$
InceptionV3 $77.1\ (72.6\text{--}81.4)$
Swin Transformer $77.4\ (72.9\text{--}81.6)$
EfficientNetV2 $78.0\ (73.7\text{--}81.9)$
MetaFormer (Base) $\mathbf{79.1}\ (\boldsymbol{74.9\text{--}83.3})$
MetaFormer-FG-2-inat-default (Ours) $\mathbf{82.2}\ (\mathbf{78.2}\textbf{--}\mathbf{86.2})$
MetaFormer-FG-2-inat-late (Ours) $\mathbf{83.6}\ (\mathbf{79.7}\textbf{--}\mathbf{87.3})$
@@ 3.1 Baseline Comparison
MetaFormer (vision-only) achieves the highest test accuracy among the compared image-only baselines (Table 3). We quantify uncertainty using 95% bootstrap confidence intervals obtained by resampling test examples with replacement (percentile method, $B=2000$). Our fusion variants, which incorporate demographic metadata, yield higher accuracy than the vision-only model under the same evaluation protocol.
3.2 Ablation Analysis
@@
We then selected the “FG-2” configuration and used the weights obtained from training it on the iNaturalist dataset [19] to create our “FG-2-inat” configuration. This surpassed all other configurations, whether using image data alone or in conjunction with metadata. This configuration reached 81.4%, a 1.5% improvement over image-only training. It seems that the fine-grained pretraining improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate.
Table 4: Joint evaluation of architecture configuration and fusion strategy reported with test accuracy (%) (95% CI).
Config Vision Late Fusion Early Fusion Default Fusion
FG-0 ${79.1}\ ({74.9\text{--}83.3})$ $79.4\ (75.1\text{--}83.3)$ $75.1\ (70.6\text{--}79.7)$ $78.5\ (74.0\text{--}82.5)$
FG-1 $77.4\ (73.2\text{--}81.6)$ $78.0\ (73.4\text{--}82.2)$ $\mathbf{76.0}\ (\mathbf{71.5}\textbf{--}\mathbf{80.2})$ $79.7\ (75.4\text{--}83.6)$
FG-2 $78.5\ (74.3\text{--}82.5)$ $80.5\ (76.3\text{--}84.5)$ $74.9\ (70.3\text{--}79.4)$ $73.4\ (68.6\text{--}78.0)$
FG-2-inat (Ours) $\mathbf{82.2}\ (\mathbf{78.2}\textbf{--}\mathbf{86.2})$ $\mathbf{83.6}\ (\mathbf{79.7}\textbf{--}\mathbf{87.3})$ $75.1\ (70.6\text{--}79.7)$ $\mathbf{82.2}\ (\mathbf{78.2}\textbf{--}\mathbf{86.2})$
Table 5: Accuracies obtained by FG-2-inat-default configuration on metadata attributes.
Attribute Accuracy (%) (95% CI)
Vision + Age $79.7\ (75.4\text{--}83.6)$
Vision + Gender $79.1\ (74.9\text{--}83.3)$
Vision + Age + Gender ${82.2}\ ({78.2}\text{--}{86.2})$
Vision + Age + Gender (late-fusion) $\mathbf{83.6}\ (\mathbf{79.7}\textbf{--}\mathbf{87.3})$
Table 6: Effect of masking schedule and probability on test accuracy.
Mask Type Mask Prob Accuracy (%) (95% CI)
Constant $0.1$ $79.7\ (75.4\text{--}83.9)$
$0.2$ $79.1\ (75.1\text{--}83.3)$
$0.3$ $79.7\ (75.1\text{--}83.6)$
$0.5$ $79.4\ (75.1\text{--}83.3)$
Linear $0.1$ $79.7\ (75.4\text{--}83.6)$
$0.2$ $78.8\ (74.6\text{--}83.1)$
$0.3$ $78.8\ (74.6\text{--}83.1)$
$0.5$ $79.4\ (75.1\text{--}83.6)$
Progressive (Ours) Scheduled $\mathbf{82.2}\ (\mathbf{78.2}\textbf{--}\mathbf{86.2})$
Too little masking encourages overreliance on metadata (hurting robustness), while too much/constant masking prevents the model from exploiting useful auxiliary cues; the default progressive schedule best balances these.
3.3 Interpretability & diagnostic performance analysis
GradCAM heatmaps (Fig. 3) demonstrate that MetaFormer attends to clinically relevant wrist regions, particularly around fracture lines and anomalous bone structures. While most heatmaps align well with pathology, some exhibit diffuse attention, suggesting occasional reliance on spurious features. This indicates that while metadata integration improves classification, further refinement is needed for precise localization. Nevertheless, the alignment of salient regions with clinical expectations strengthens confidence in the model’s decision process. Table 7 shows per-class sensitivity, specificity, & precision scores for configuration “FG-2-inat”. Similarly, as shown in Fig. 4, metadata enhances fracture detection while preserving performance in bone anomaly and soft tissue recognition.
Table 7: Evaluation of our default fusion (FG-2-inat).
Configuration Cls Sensitivity (%) Specificity (%) Precision (%)
Vision Only 0 $62.2$ $98.7$ $96.1$
11 $100.0$ $82.9$ $75.0$
22 $77.4$ $88.3$ $76.1$
Vision+Meta 0 66.4{66.4} $98.7$ 96.3{96.3}
11 $100.0$ $82.9$ $75.0$
22 $77.4$ 90.4{90.4} 79.5{79.5}
Table 8: Performance evaluation on a significantly larger (fracture vs. no fracture) dataset.
FG-2-inat (Ours) Test Accuracy (%)
Vision Only $50.1$
Vision + Meta $\mathbf{60.4}$
@@ 3.3 Interpretability & diagnostic performance analysis
3.4 Failure cases & error analysis
[…]
@@ 4 Conclusion
Overall, our experiments demonstrate three key findings: (1) Fine-grained transformer architecture outperforms conventional CNNs for pediatric wrist pathology. (2) Demographic metadata (age, sex) significantly improves recognition across both limited and large-scale datasets, with late fusion being most effective. (3) Fine-grained pretraining enhances generalization compared to using ImageNet weights. These results emphasize that integrating developmental context into vision models is essential for robust pediatric musculoskeletal diagnosis.